OpenStudy (anonymous):

Find the angle theta to the nearest minute for sin theta=0.1655. a. 9 degrees 32' b. 9 degrees 38' c. 9 degrees 39' d. 9 degrees 31'

OpenStudy (anonymous):

\[\sin(\theta) = 0.1655 \implies \theta = \sin^{-1}(0.1655)\]

OpenStudy (anonymous):

ans= (d) 9 degrees 31'

OpenStudy (anonymous):

\[\large \theta = 9.5262\] Clearly 9 degrees we have but we have to find minute also.. 1 hour = 60 minutes: \[\large 5262 \times 60 = 315720\] So: 9 degrees and 31 minutes..

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Welcome dear..